OpenStudy (anonymous):

A stunt man drives a car at a speed of 20 m/s off a 20-m-high cliff. The road leading to the cliff is inclined upward at an angle of 20∘.

OpenStudy (lieutenantgeneral):

car O(launch speed (u) at angle 20 with ox o____________ x |.... |(h).... |............. |................. O lands in curve parabola vertical speed (down) = u sin 20 hori speed = u cos 20 ....... (same always) y = u sin 20 *t + 0.5 g t^2 = 30 x = u cos 20 * t = R (range) here (t) time of flight > eliminate 20 sin 20 *t + 0.5*9.8 t^2 = 30 6.84 t + 4.9 t^2 - 30 = 0 it gives >> t = + 1.873 s (leaving other -ve time) R = range where lands = 20* cos 20 * 1.873 R = 35.2 meter from base ============================ impact speed energy conservation> initial = final 0.5 mu^2 + mgh = 0.5 mv^2 + 0 (at base) u^2 + 2gh = v^2 v^2 = 20*20 + 2*9.8*30 = 400 + 588 = 988 v = impact speed = 31.43 m/s other method>>> vy = dy/dt = u sin 20 + g t = 25.195 vx = u cos 20 = 18.794 v^2 = [vx^2+ vy^2]^1/2 = 31.43 m/s

OpenStudy (anonymous):

How far from the base of the cliff does the car land?